# Masked Mixers for Language Generation and Retrieval

### v2

@@ 4.1 Architecture Optimizations
We perform extensive architectural experiments to optimize the masked mixer’s architecture using TinyStories, and found that the use of input representation is an effective guide to training efficiencies for this model. Most notably, we modify the original MLP Mixer architecture to use only one convolution and no nonlinear activation between sequence elements as this was found to exhibit superior input representation accuracy and more efficient training (c.f. Section 9.1). The number of layers, hidden layer width, batch sizes, and hyperparameters like learning rates (all with linear decay) were then optimized on the same dataset via independent line searches. We found that multi-headed masked mixers and masked mixers with non-unitary convolutional kernel sizes were no more efficient, and Softmax-transformed convolutional weight mixers were substantially less efficient, than the standard one-masked-convolution architecture for TinyStories CLM training (Section 9.2).
@@
We found that the optimized masked mixer exhibited more efficient TinyStories training characteristics than non-optimized Llama-style transformers (Section 9.3) but considered this to be an unfair comparison and proceeded to optimize the transformer model. Most notably here, we found that the default self-attention head number was much too large for optimal training on this dataset and reduced the default (32 heads) to four (Section 9.4). As for the masked mixer, we also optimized the number of hidden layers, layer widths, batch sizes, and learning rates using the same dataset. Once this was performed, it was found that the optimized Llama-style transformer was somewhat more efficient for CLM training on TinyStories than the optimized masked mixers.
@@
As a general note, most comparisons between masked mixers and transformers in this work are performed between models with the same number of layers (usually $n_{l}=8$ or $n_{l}=16$) but with $d_{m}=512$ layer width for transformers and $d_{m}=1024$ for masked mixers. This difference in $d_{m}$ is the result of the optimizations detailed above: masked mixers require much less device memory as they contain fewer inter-token parameters (Tables S8 and S9), and as such these models are approximately equivalent with respect to on-device memory during training. During the aforementioned optimizations, we find empirically that increasing the transformer’s width to $d_{m}=1024$ with a decrease in batch size results in worse training efficiency and likewise decreasing the masked mixer’s width to $d_{m}=512$ while increasing the batch size leads to worse efficiency. We consider this model-dependent optimization as vital to valid CLM efficiency comparisons, as otherwise one can simply choose architectural or hyperparameter constraints that favor one or the other model. We open our optimization results open to the public in this work and urge other researchers to do the same, as the failure to perform at least some reasonable amount of architectural or hyperparameter optimization leads to relatively meaningless comparisons between architectures.
@@ 4.2 Autoregressive inference for masked mixers
We introduce a simple inference method that makes use of all the information that the masked mixers learns per token position during training. Recall that the mixer has effectively a fixed positional encoding intrinsic in the model’s convolutional weights, and that these weights are responsible for all of the inter-token information flow. One can make use of this positional encoding by iterating through the context indices while adding the token generated from the mixer’s output at each index.
A Python implementation of this approach is as follows:
@@
whereas a transformer with a slightly higher loss yields the the following completion with somewhat more deviation from the plot.
@@
To test the ability of masked mixers to be trained efficiently for causal language modeling on datasets that better represent those used for large frontier models while maintaining a reasonable training time on our limited compute we turned to a 10 billion token subset of the FineWeb-edu [24]. We train a Llama-3 [5] tokenizer of size 8000 on this dataset before using this tokenizer to train randomly-initialized transformers and mixers.
We first profiled the Llama transformer model to determine the optimal number of attention heads for this larger dataset, and found that, as for TinyStories, a 4-headed model achieved the lowest per-step loss and required the least time per step (Figure S7). We next trained transformers and masked mixers on the FineWeb dataset with varying context windows and batch sizes according to the rule $n_{ctx}*n_{b}=65536$ to prevent memory overflow, and the results of these experiments are shown in Figure 4. We observe that as the language model’s mapping goes from $\mathbb{N}^{1024}\to\mathbb{N}$ to $\mathbb{N}^{128}\to\mathbb{N}$ (and becomes closer to a bijective and thus invertible map) the masked mixer’s relative performance increases compared to the transformer.
@@ 4.4 Masked mixers are more efficient language learners than early transformer implementations
In some sense it is unsurprising that current transformer implementations are somewhat more efficient learners for large context windows than masked mixers because transformers have seen a number of architectural improvements and compute optimizations since their introduction. It is useful therefore to compare the learning efficiencies of early transformer implementations of all-next-token causal language modeling to masked mixers with the same training method, as these can be thought of as being similar in their developmental stage. We chose a model that was introduced before improvements such as Rotary Positional Encoding [26] and Flash Attention 2 and sophisticated parameter initialization techniques to name a few.
@@ 5.1 Neither Dataset nor Task Stochasticity do not explain mixer transformer performance differences
One explanation of the improved performance of masked mixers on smaller context windows and the improved performance of transformers on larger context windows is that transformers are better able to filter out noise via attention, and this noise could exist in the dataset itself or else in the language modeling task. We first investigated whether the gaps between transformer and masked mixer causal language model training efficiency were due to stochasticity inherent in the dataset, by which we mean the somewhat random structure inherent in natural languages. For example, observe that many equally valid next words could be used for most of the words in most English sentences such that any one of these could be picked at random, as nouns and verbs and adjectives are often free interchanged and grammars are relatively inconsistent between passages.
On straightforward way to test this is by observing CLM loss for each model type during training on a dataset that has a higher proportion of formal language text with fewer degrees of freedom. We chose to train on mathematical text, the FineMath 4+ dataset with approximately 10 billion tokens [27], as an example of such a dataset. After CLM training we find that lower cross-entropy loss is achieved as expected for a more deterministic dataset, but that once again transformers are more efficient for longer-context but not shorter-context training (Figure S6), supporting the hypothesis that inherent dataset stochasticity does not contribute to the transformer-mixer training efficiency gap.
@@
As shown in Figure 6, we see that there is an almost identical loss gap between masked mixer and transformer for bidirectional training compared to left-token CLM training, providing evidence for the idea that task stochasticity does not explain the relative performance of each model type. Together with the earlier results in this section, we conclude that neither dataset nor task-specific stochasticity explain the mixer versus transformer training efficiency differences.
@@ 5.2 Multi-token and many-token prediction efficiencies
After observing that masked mixers perform worse relative to transformers as $n$ increases during CLM training in which models learn $\mathbb{N}^{n}\to\mathbb{N}^{1}$ mappings, we investigated the relative training efficiencies of these models on language modeling tasks described as multifunctions of $\mathbb{N}^{n}\to\mathbb{N}^{m},m>1$ and more specifically $\mathbb{N}^{512}\to\mathbb{N}^{2}$.
The assumptions behind training a language model to predict one next token at a time is that this is an easier task than predicting many next tokens at once, and that simply being able to predict each next token allows for effective inference. The latter implies that a greedy strategy during training is best, but the ability of inference methods such as beam search to outperform greedy inference suggests that the training a model to predict multiple ‘next’ tokens may be beneficial. This has recently been referred under various names such as ‘multiple token prediction’ or ‘non-myopic pretraining’, and there are many method by which one could implement this idea ranging from a multi-headed approach where each head predicts one token [29] to the more direct approach of performing $n$ forward passes and add the corresponding loss values for each of $n$ tokens predicted, performing label shifting with a shift position size corresponding to the number of forward passes similar to that in [30]. We take the latter approach and perform two forward passes through the model for two-token prediction. As shown in Figure 7 (a), the loss gap between masked mixer and transformer is somewhat smaller for two-token prediction compared to one-token prediction at 1.8% and 2.0%, respectively (refer to Figure 4).
We also investigated the efficiencies of language models trained to predict many next tokens in one forward pass instead of only one such that each token prediction is $\mathbb{N}^{n/2}\to\mathbb{N}$ with all token predictions being $\mathbb{N}^{n/2}\to\mathbb{N}^{n/2}$. Somewhat arbitrarily choosing to generate half of the text (meaning a context window of 512 we use the first 256 tokens as inputs and perform loss computation on the next 256 tokens), we find that a masked mixer and transformer yield nearly identical Fineweb training efficiencies Figure 7 (c).
@@ 5.3 Masked Mixers are far better Autoencoders than Transformers
At the other extreme of the $\mathbb{N}^{n}\to\mathbb{N}^{1}$ mapping that is performed during next token prediction, we test masked mixer and transformer training efficiencies on a task that requires an $\mathbb{N}^{n}\to\mathbb{N}^{n}$ (i.e. invertible) transformation: a non-trivial (with respect to inter-token information flow) autoencoding in which an encoder takes the input and compresses it to one hidden layer, and the decoder takes this hidden layer and attempts to regenerate the input tokens. Such an autoencoder may be implemented by using a encoder-decoder model where a standard masked mixer acts as the encoder, the last hidden state of the last token is taken as the embedding, and a decoder composed of a stack of masked mixer modules transforms this embedding (repeated for each input word) into a sequence to match the input as shown in 8. We implement an analogous architecture using a modified Llama-style transformer without word-token embedding layers or language modeling heads for encoder and decoder modules.
This mixer autoencoder reaches a cross-entropy loss on FineMath comparable to non-optimized transformers and mixers (albeit with around twice the number of training steps required) and very little or no overfitting, indicating that the embedding created by the model’s encoder provides generalizable information on the input to the decoder, and in particular enough information for the decoder to reconstruct the entire input sequence approximately as accurately as trained all-next-token causal language models. In contrast, transformers are far less efficient for this autoencoding and are unable to predict input tokens with anywhere near the accuracy of causal language modeling-trained transformers or masked mixers or mixer autoencoders (Figure 8). We observe this to be the case for TinyStories as well, and confirm that these differences are not caused by differences in $d_{m}$ between transformer and masked mixer (Figure S13).
@@ 6 Masked Mixers are Better for Retrieval
Along with generation one of the most frequently encountered language tasks today is retrieval, a task which usually involves finding one or more matches for a given language segment given a set of text embeddings. One of the most common uses for this is in retrieval-augmented generative search, whereby a search phrase is ‘read’ by a language model and the resulting embedding from that phrase is then used to match one or more segments of a corpus of text, which is in turn fed to a generative model along with the original search phrase [31].
As is the case for language generation, nearly every retrieval model used today is based on the transformer architecture. The results in this work lead to the hypothesis that attention is not well-suited for the task of retrieval because these transformations are biased towards non-invertibility, such that a mapping with this transformation becomes many-to-one (Figure 3). For predicting a single next token it appears that this mapping is not detrimental and may even be beneficial, as masked mixers learn to reduce the information passing to deep layers during causal language model training.
@@ 6.1 Synthetic dataset generation and embedding
To investigate retrieval abilities of transformers and masked mixers on in-distribution data while maintaining flexibility as to the number of samples included per retrieval, we generated a synthetic summary sentence for the first 200k samples (truncated to 512 tokens maximum) in the Finewen dataset using Llama-3 (8b) Instruct [5] by prompting that model to read each 512-token-length text segment and give a one-sentence summary of that segment. We found that this relatively small model was more than capable of accurately and concisely summarizing each segment. We repeated this procedure with the FineMath 4+ dataset, for either 200k or 400k samples, and changed the model prompt to make the summary query more concise (and more difficult to match).
Each text segment and its corresponding summary were fed to a pretrained Llama-style transformer or masked mixer and the last hidden layer activations were saved as the model’s embedding. We use the second-to-last token’s output for the embedding because the ’last’ masked mixer token is untrained due to the index shift that occurs during all-next-token training of these models. This procedure only rarely would be expected to lead to input information loss, as the synthetic summary queries nearly all contain ending punctuation tokens. As masked mixers have an absolute positional encoding, we pretrain these models on left-padded tokens rather than the right-padded token segments used elsewhere in this work in order to be able to use the second-to-last token hidden layer for inputs of all sizes. We experimentally determined that left versus right padding during pretraining yields virtually no difference for subsequent retrieval training for transformers, as long as the retrieval training itself proceeds with left padding. For consistency, we pretrained transformers used for retrieval with left-padded inputs as well.
@@ 6.2 Indirect Retrieval Model Architecture and Training Approach
For our retrieval model we chose to implement a mixer with bidirectional convolutions rather than masked ones, similar to MLP mixers but with only one convolutional layer rather than two between sequence elements. This was chosen because a comparison of all possible matches requires information from all input elements by definition, making transformer models unsuitable. This was experimentally verified, as transformers were found to be very poor retrieval models even for small comparison contexts (Table S7). Word -token embedding layers are removed (as the inputs are embeddings already) as are language heads, and instead each last hidden layer has a $d_{model}\to 1$ transformation where the unitary output corresponds to the likelihood (after Softmax transformation) of that input embedding being the correct match. Softmax transformation serves to stabilize the retrieval model’s loss, as increasing all logit values does not decrease the cross-entropy loss, and a depiction of this architecture for masked mixers is found in Figure 9. We make use of the Pytorch Cross-Entropy Loss function’s compatibility with class probabilities for this model.
Arbitrarily choosing the first ‘token’ to be the summary and all other tokens to be the potential matches, each input is then assembled by randomly sampling all story embeddings before replacing one input (at a random location) with the embedding of the matching story. A non-optimized version of this approach for generating inputs for the retrieval model is found in Algorithm 1.
@@
1:$x=O_{l}(\mathrm{summaries},\theta_{g})$triangleright Summary embeddings
2:$y=O_{l}(\mathrm{stories},\theta_{g})$triangleright Story embeddings
@@
4:vector a: len(a) = c $\triangleright$ Vector of embeddings for model input
@@
7:$n\leftarrow i:0<=i<N$
8:while n $\leftarrow$ 0, n++, n < N do $\triangleright$ Loop over summary/story pairs
9:     a[0] $\leftarrow$ x[n] $\triangleright$ First input element is the summary to match
10:     weights: weights[0, 1, … N] $\leftarrow$ 1
11:     weights[n] $\leftarrow$ 0
12:     r = multinomial(weights, c-1) $\triangleright$ Random sample of all input indices except matching
13:     a[1:c] $\leftarrow$ y[r] $\triangleright$ Replace elements of a with lookup of embeddings
14:     m = randint(1, c) $\triangleright$ Random index to place matching story embedding
15:     a[m] $\leftarrow$ y[n] $\triangleright$ Matching story embedding placed
16:     q[m] $\leftarrow$ n $\triangleright$ One-hot label at position m
17:     yield a, q
18:end while
The outputs after Softmax transformation are compared to the true distribution via standard cross-entropy loss, which is then back-propagated through the retrieval model but not further to the generative model or input embedding. Note that the logits are not Softmax-transformed during inference as this operation has no effect on top-k choice without noise introduction. We use a relatively modest batch size (512) compared to most retrieval model training approaches. In contrast to the rest of this work, the efficiency of an embedding for retrieval modeling is not measured in loss per compute amount or even loss per number of samples seen but instead minimum evaluation cross-entropy loss over a pre-determined.
@@
Figure 9: Retrieval mixer architecture and FineWeb retrieval evaluation cross-entropy loss (sampled without replacement).

### v3

@@ 4.1 Architecture optimizations
We perform extensive experiments to optimize the masked mixer’s architecture using TinyStories, and find that the use of input representation is an effective guide to training efficiencies for this model. Most notably, we modify the original MLP Mixer architecture to use only one convolution and no nonlinear activation between sequence elements as this was found to exhibit superior input representation accuracy and more efficient training (cf. Section 9.1). The number of layers, hidden layer width, batch sizes, and hyperparameters like learning rates were then optimized on the same dataset via independent line searches. We found that multi-headed masked mixers and masked mixers with non-unitary convolutional kernel sizes were no more efficient, and Softmax-transformed convolutional weight mixers were substantially less efficient, than the standard one-masked-convolution architecture for TinyStories CLM training (Section 9.2).
@@
We found that the optimized masked mixer exhibited more efficient TinyStories training characteristics than non-optimized Llama-style transformers (Section 9.3) but considered this to be an unfair comparison and proceeded to optimize the transformer model’s architecture. We found that the default self-attention head number was much too large for optimal training on this dataset and reduced the default (32 heads) to four (Section 9.4). As for the masked mixer, we also optimized the number of hidden layers, layer widths, batch sizes, and learning rates using the same dataset. Once this was performed, it was found that the optimized Llama-style transformer was somewhat more efficient for CLM training on TinyStories than the optimized masked mixers.
@@
As a general note, most comparisons between masked mixers and transformers in this work are performed between models with the same number of layers (usually $n_{l}=8$ or $n_{l}=16$) but with $d_{m}=512$ layer width for transformers and $d_{m}=1024$ for masked mixers. This difference in $d_{m}$ is the result of the optimizations detailed above: masked mixers require much less device memory as they contain fewer inter-token parameters (Tables S9 and S10), and as such these models are approximately equivalent with respect to on-device memory during training. During the aforementioned optimizations, we find empirically that increasing the transformer’s width to $d_{m}=1024$ with a decrease in batch size results in worse training efficiency and likewise decreasing the masked mixer’s width to $d_{m}=512$ while increasing the batch size leads to worse efficiency. We consider this model-dependent optimization as vital to valid CLM efficiency comparisons, as otherwise one can simply choose architectural or hyperparameter constraints that favor one or the other model. We open our optimization results open to the public in this work and urge other researchers to do the same, as the failure to perform at least some reasonable amount of architectural or hyperparameter optimization leads to relatively meaningless efficiency comparisons between architectures.
@@ 4.2 Autoregressive inference for masked mixers
We introduce a simple inference method that makes use of all the information that the masked mixers learns per token position during training. Recall that the mixer has effectively a fixed positional encoding intrinsic in the model’s convolutional weights, and that these weights are responsible for all of the inter-token information flow. One can make use of this positional encoding by iterating through the context indices while adding the token generated from the mixer’s output at each index. A Python implementation of this approach is as follows:
@@
whereas a transformer with a slightly higher loss gives the following:
@@
To test the ability of masked mixers to be trained efficiently for causal language modeling on datasets that better represent those used for large frontier models while maintaining a reasonable training time on our limited compute, we turned to a 10 billion token subset of the FineWeb-edu [24]. We train a Llama-3 [5] tokenizer of size 8000 on this dataset before using this tokenizer to train randomly-initialized transformers and mixers.
We first profiled the Llama transformer model to determine the optimal number of attention heads for this larger dataset, and found that, as for TinyStories, a 4-headed model achieved the lowest per-step loss and required the least time per step (Figure S10). We next trained transformers and masked mixers on the FineWeb dataset with varying context windows and batch sizes according to the rule $n_{ctx}*n_{b}=65536$ to prevent memory overflow, and the results of these experiments are shown in Figure 4. We observe that as the language model’s mapping goes from $\mathbb{N}^{1024}\to\mathbb{N}$ to $\mathbb{N}^{128}\to\mathbb{N}$ (and becomes closer to a bijective and thus invertible map) the masked mixer’s relative training efficiency increases compared to the transformer.
@@ 4.4 Masked mixers are more efficient language learners than early transformer implementations
In some sense it is unsurprising that current transformer implementations are somewhat more efficient learners for large context windows than masked mixers because transformers have seen a number of architectural improvements and compute optimizations since their introduction. It is useful therefore to compare the learning efficiencies of early transformer implementations of all-next-token causal language modeling to masked mixers with the same training method, as these can be thought of as being similar in their developmental stage. We chose a model that was introduced before improvements such as Rotary Positional Encoding [26] and Flash Attention 2 [27] and sophisticated parameter initialization techniques to name a few.
@@ 5.1 Neither dataset nor task stochasticity do not explain mixer transformer performance differences
One explanation of the decrease in relative training efficiency of masked mixers to transformers as the context window increases is that transformers are better able to filter out noise via attention, where this noise could exist in the dataset itself or else in the language modeling task. We first investigated whether the gaps between transformer and masked mixer causal language model training efficiency were due to stochasticity inherent in the dataset, by which we mean the somewhat random structure inherent in natural languages. For example, observe that many equally valid next words could be used for most English sentences such that any one of these could be picked at random, as nouns and verbs and adjectives are often free interchanged and grammars are relatively inconsistent between passages.
One straightforward way to test this is by observing CLM losses for each model type during training on a dataset that has a higher proportion of formal language to natural language. We chose to train on mathematical text, the FineMath 4+ dataset with approximately 10 billion tokens [28], as an example of such a dataset. After CLM training we find that lower cross-entropy loss is achieved as expected for a more deterministic dataset, but that once again transformers are more efficient for longer-context but not shorter-context training (Figure S9), supporting the hypothesis that inherent dataset stochasticity does not contribute to the transformer-mixer training efficiency gap.
@@
As shown in Figure 6, there is an almost identical loss gap between masked mixer and transformer for bidirectional training compared to left-token CLM training, providing evidence for the idea that task stochasticity does not explain the relative performance of each model type.
@@ 5.2 Multi-token and many-token prediction efficiencies
After observing that masked mixers perform worse relative to transformers as $n$ increases during CLM training in which models learn $\mathbb{N}^{n}\to\mathbb{N}^{1}$ mappings, we investigated the relative training efficiencies of these models on language modeling tasks described as multifunctions of $\mathbb{N}^{n}\to\mathbb{N}^{m},m>1$ and more specifically we start by investigating a language that may be described as a mapping $\mathbb{N}^{512}\to\mathbb{N}^{2}$.
The assumptions behind training a language model to predict one next token at a time is that this is an easier task than predicting many next tokens at once, and that simply being able to predict each next token allows for effective inference. The latter implies that a greedy strategy during training is best, but the ability of inference methods such as beam search to outperform greedy inference suggests that the training a model to predict multiple ‘next’ tokens may be beneficial. This has recently been referred under various names such as ‘multiple token prediction’ or ‘non-myopic pretraining’, and there are many method by which one could implement this idea ranging from a multi-headed approach where each head predicts one token [30] to the more direct approach of performing $m$ forward passes and add the corresponding loss values for each of $m$ tokens predicted, performing label shifting with a shift position size corresponding to the number of forward passes similar to the method used in [31]. We take the latter approach and perform two forward passes through the model for two-token prediction. As shown in Figure 7 (a), the loss gap between masked mixer and transformer is slightly smaller for two-token prediction compared to one-token prediction at 1.8% and 2.0%, respectively (refer to Figure 4).
We also investigated the efficiencies of language models trained to predict many next tokens in one forward pass instead of only one such that each token prediction is $\mathbb{N}^{n/2}\to\mathbb{N}^{1}$ with all token predictions being $\mathbb{N}^{n/2}\to\mathbb{N}^{n/2}$ which was achieved by the architecture shown in Figure 7 (b). Somewhat arbitrarily choosing to generate half of the text (meaning a context window of 512 we use the first 256 tokens as inputs and perform loss computation on the next 256 tokens), we find that a masked mixer and transformer yield nearly identical Fineweb training efficiencies Figure 7 (c).
@@ 5.3 Masked mixers are far better autoencoders than transformers
At the other extreme of the $\mathbb{N}^{n}\to\mathbb{N}^{1}$ mapping that is performed during next token prediction, we test masked mixer and transformer training efficiencies on a task that requires an $\mathbb{N}^{n}\to\mathbb{N}^{n}$ (i.e. invertible) transformation: a non-trivial (with respect to inter-token information flow) autoencoding in which an encoder takes the input and compresses it to one hidden layer, and the decoder takes this hidden layer and attempts to regenerate the input tokens. Such an autoencoder may be implemented by using a encoder-decoder model where a standard masked mixer acts as the encoder, the last hidden state of the last token is taken as the embedding, and a decoder composed of a stack of masked mixer modules transforms this embedding (repeated for each input word) into a sequence to match the input as shown in Figure 8. We implement an analogous architecture using a modified Llama-style transformer without word-token embedding layers or language modeling heads for encoder and decoder modules.
This mixer autoencoder reaches a cross-entropy loss on FineMath comparable to non-optimized transformers and mixers (albeit with around twice the number of training steps required) and very little or no overfitting, indicating that the embedding created by the model’s encoder provides generalizable information on the input to the decoder, and in particular enough information for the decoder to reconstruct the entire input sequence with accuracies approaching all-next-token causal language models trained using identical compute (Figure 8). In contrast, transformers are far less efficient for this autoencoding and are unable to predict input tokens with anywhere near the accuracy of causal language modeling-trained transformers or masked mixers or mixer autoencoders (Figure 8). We observe this to be the case for TinyStories as well, and confirm that these differences are not caused by differences in $d_{m}$ between transformer and masked mixer (Figure S13).
@@ 6 Masked Mixers are Better for Retrieval
One of the most frequently encountered language tasks today is retrieval, a task which usually involves finding one or more matches (the ‘target’) for a given language segment (the ‘query’) given a set of potential matches. One of the most common uses for this is in retrieval-augmented generative search, whereby a search phrase is ‘read’ by a language model and the resulting embedding from that phrase is then used to match one or more segments of a corpus of text, which is in turn fed to a generative model along with the original search phrase [32].
As is the case for language generation, nearly every retrieval model used today is based on the transformer architecture. The results in previous sections of this work lead to the hypothesis that attention is not well-suited for the task of retrieval because these transformations are biased towards non-invertibility, such that a mapping with this transformation becomes many-to-one (Figure 3). For predicting a single next token it appears that this mapping is not detrimental and may even be beneficial, as masked mixers learn to reduce the information passing to deep layers during causal language model training.
@@ 6.1 Synthetic dataset generation and embedding
To investigate retrieval abilities of transformers and masked mixers on in-distribution data while maintaining flexibility as to the number of samples included per retrieval, we generated a synthetic summary sentence for the first 200k samples (truncated to 512 tokens maximum) in the FineWeb dataset using Llama-3 (8b) Instruct [5] by prompting that model to read each 512-token-length text segment and give a one-sentence summary of that segment. We found that this relatively small model was more than capable of accurately and concisely summarizing most target segments. We repeated this procedure with the FineMath 4+ dataset, for either 200k or 400k samples, and changed the model prompt to make the summary query more concise (and more difficult to match).
Each text segment and its corresponding summary were fed to a pretrained Llama-style transformer or masked mixer and the last hidden layer activations were saved as the model’s embedding. We use the second-to-last token’s output for the embedding because the ‘last’ masked mixer token is untrained due to the index shift that occurs during all-next-token training of these models. This procedure only rarely would be expected to lead to input information loss, as the synthetic summary queries nearly all contain ending punctuation tokens. As masked mixers have an implicitly defined absolute positional encoding, we pretrain these models on left-padded tokens rather than the right-padded token segments used elsewhere in this work in order to be able to use the second-to-last token hidden layer for inputs of all sizes. Training with left-padded tokens leads to some decrease in CLM training efficiency as no weight masking is performed on these inputs. We experimentally determined that left versus right padding during pretraining yields virtually no difference for subsequent retrieval training for transformers, as long as the retrieval training itself proceeds with left padding. For consistency, we pretrained transformers used for retrieval with left-padded inputs as well.
@@ 6.2 Indirect retrieval model architecture and training approach
For our retrieval model we chose to implement a mixer with bidirectional convolutions rather than masked ones, similar to MLP mixers but with only one convolutional layer rather than two between sequence elements, as transformers were found to be very poor retrieval models even for small comparison contexts (Table S7). Word -token embedding layers are removed (as the inputs are embeddings already) as are language heads, and instead each last hidden layer has a $d_{model}\to 1$ transformation where the unitary output corresponds to the likelihood (after Softmax transformation) of that input embedding being the correct match. Softmax transformation serves to stabilize the retrieval model’s loss, as increasing all logit values does not decrease the cross-entropy loss, and a depiction of this architecture for masked mixers is found in Figure 9.
Arbitrarily choosing the first ‘token’ to be the summary and all other tokens to be the potential matching target text segment embeddings, each input is then assembled by randomly sampling all target embeddings before replacing one input (at a random location) with the embedding of the matching target. A non-optimized version of this approach for generating inputs for the retrieval model is found in Algorithm 1.
@@
1:$x=O_{l}(\mathrm{queries},\theta_{g})$triangleright Query embeddings
2:$y=O_{l}(\mathrm{targets},\theta_{g})$triangleright Target embeddings
@@
4:vector a: len(a) = c $\triangleright$ Vector of embedding vectors for model input
@@
7:while n $\leftarrow$ 0, n++, n < N do $\triangleright$ Loop over query/target pairs
8:     a[0] $\leftarrow$ x[n] $\triangleright$ First input element is the query to match
9:     weights: weights[0, 1, … N] $\leftarrow$ 1
10:     weights[n] $\leftarrow$ 0
11:     r = multinomial(weights, c-1) $\triangleright$ Random sample of all input indices except matching
12:     a[1:c] $\leftarrow$ y[r] $\triangleright$ Replace elements of a with lookup of embeddings
13:     m = randint(1, c) $\triangleright$ Random index to place matching target embedding
14:     a[m] $\leftarrow$ y[n] $\triangleright$ Matching target embedding placed
15:     q[m] $\leftarrow$ n $\triangleright$ One-hot label at position m
16:     yield a, q, m
17:end while
The outputs after Softmax transformation are compared to the true distribution via standard cross-entropy loss, which is then back-propagated through the retrieval model but not further to the generative model or input embedding. Note that the logits are not Softmax-transformed during inference as this operation has no effect on top-k choice without noise introduction. We use a relatively modest batch size (512). We make use of the Pytorch Cross-Entropy Loss function’s compatibility with class probabilities for this model. In contrast to the rest of this work, the efficiency of an embedding for retrieval modeling is not measured in loss per compute amount or even loss per number of samples seen but instead minimum evaluation cross-entropy loss over a pre-determined.
@@
Figure 9: Retrieval mixer architecture and FineWeb retrieval evaluation cross-entropy loss ($n$ samples per retrieval model forward pass, sampled without replacement).